a vessel contains a non linear triatomic gas.if 50% of gas dissociate into individual atom, then find new value of freedom by ignoring the vibrational mode and any further dissociation

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
Let's assume we have1mole of triatomic gas∴3Nais present
So,0.5moles=1.5Naatoms
1part of0.5moles remains untouched
Degree of dissociation=0.5×6=3
Degree of freedom for0.5Na=1.5×0.5=0.75
Total=3+0.75=3.75